OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. A. 84 B. 7 C. -12

OpenStudy (anonymous):

ratio of y/x = 7/1 y = 7x => x = -12 when y = -84 just divide y by 7

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (anonymous):

no pb :)